# Spectropolarimetric Inversion in Four Dimensions with Deep Learning (SPIn4D):

### v1

@@ Abstract
The National Science Foundation’s Daniel K. Inouye Solar Telescope (DKIST) will provide high-resolution, multi-line spectropolarimetric observations that are poised to revolutionize our understanding of the Sun. Given the massive data volume, novel inference techniques are required to unlock its full potential. Here, we provide an overview of our “SPIn4D” project, which aims to develop deep convolutional neural networks (CNNs) for estimating the physical properties of the solar photosphere from DKIST spectropolarimetric observations. We describe the magnetohydrodynamic (MHD) modeling and the Stokes profile synthesis pipeline that produce the simulated output and input data, respectively. These data will be used to train a set of CNNs that can rapidly infer the four-dimensional MHD state vectors by exploiting the spatiotemporally coherent patterns in the Stokes profile time series. Specifically, our radiative MHD model simulates the small-scale dynamo actions that are prevalent in quiet-Sun and plage regions. Six cases with different mean magnetic fields have been conducted; each case covers six solar-hours, totaling TB in data volume. The simulation domain covers at least Mm with km spatial resolution, extending from the upper convection zone up to the temperature minimum region. The outputs are stored at a 40 s cadence. We forward model the Stokes profile of two sets of Fe I lines at and nm, which will be simultaneously observed by DKIST and can better constrain the parameter variations along the line of sight. The MHD model output and the synthetic Stokes profiles are publicly available.
UTF8gbsn
@@ 1 Introduction
With the help of the above new facilities, spectropolarimetric observations (typically in the form of wavelength dependent Stokes profiles) for solar physics will significantly expand in the near future. These advanced observations are sensitive to the plasma state variables across the solar atmosphere (del Toro Iniesta & Ruiz Cobo, 2016; Bellot Rubio & Orozco Suárez, 2019, and reference therein), offering a unique tool to probe the solar environment in different heights through sophisticated inversion techniques, whereby the multi-dimensional state of the plasma is inferred from 2D maps of the polarized spectra (Ruiz Cobo & del Toro Iniesta, 1992; Asensio Ramos et al., 2008; Socas-Navarro et al., 2015; Milić & van Noort, 2018; delaCruzRodriguez et al., 2019; Quintero Noda et al., 2021, 2023; Ruiz Cobo et al., 2022; Asensio Ramos & Díaz Baso, 2019; Pastor Yabar et al., 2019; Borrero et al., 2019). For example, initial DKIST/ViSP data measurements of magnetic fields in both quiet and plage regions highlight the utility of spectropolarimetric observation in probing plasma dynamics in otherwise inaccessible environments (Campbell et al., 2023; da Silva Santos et al., 2023; Kuridze et al., 2024). Developing good inversion techniques is thus required to tackle a host of unanswered questions in both solar and plasma physics in general, for instance, probing the mechanisms behind local magnetic dynamos (Vögler & Schüssler, 2007; Stenflo, 2012; Rempel, 2014; Lord et al., 2014), understanding the flux of helicity and energy across the solar surface (Welsch, 2015; Schuck & Antiochos, 2019; Lumme et al., 2019; Thalmann et al., 2021; Liu & Schuck, 2012; Liu et al., 2014, 2023), and unraveling the magnetic foundations necessary for solar eruptions and coronal heating (Antiochos, 1998; Antiochos et al., 1999; Moore et al., 2001; Priest & Forbes, 2002; Sun et al., 2013; Wang et al., 2015; Liu et al., 2016; Chitta et al., 2017; Wyper et al., 2017; Wang et al., 2017; Samanta et al., 2019).
@@
To address the challenges outlined above and take advantage of the potential of DL, we launched the “Spectropolarimetric Inversion in Four Dimensions with Deep Learning” (SPIn4D) project222https://ifauh.github.io/SPIN4D/ to train CNN models on radiative MHD (RMHD) simulations with a significantly larger dataset compared to previous efforts. These data are publicly available333http://dtn-itc.ifa.hawaii.edu/spin4d/DR1/. Figure 2 presents a schematic overview of key elements of each phase of the project. Our focus will be on photospheric regions with intermediate field strengths (between quiet sun and plage, or up to a few hundred Gauss when spatially averaged), characterized by relatively simple field geometries. These regions are expected to be prevalent in the initial years of DKIST’s operation. Simulations are carried out using the Max-Planck University-of-Chicago Radiative MHD code (MURaM, Vögler et al., 2005; Rempel et al., 2009; Cheung et al., 2010; Rempel, 2012, 2014) and take the quiet Sun small-scale dynamo simulation (SSD) of Rempel (2014) as the point of departure. Next, we employ forward-modeling of radiative transfer through these simulations using a new version of the SIR code (SIR3D) and the Departure Coefficient aided Stokes Inversion based on Response Functions (DeSIRe, Ruiz Cobo et al., 2022) code to synthesize Stokes profiles for multi-line observations. We have selected the well-studied Fe I lines at nm, nm, nm, and nm due to their significant Landé factors (, , , and , respectively) and their wide-spread use in ground and space-based observations with high Stokes SNR. The third phase involves developing CNN models that aim to accurately correlate time series of observational data, especially those from the DL-NIRSP instrument at DKIST, with precise 4D MHD states. The models will be rigorously trained and evaluated using the data generated in the first two steps. The ability of this model will be compared with the SIR inversion code as a base line. Higher-level variables such the vector velocity and associated Poynting flux across the photosphere may also be estimated directly. The comprehensive workflow and expected results are outlined in Figure 2.
@@ 2 Solar Atmosphere Simulation
All Cases 1–6 were simulated for about six physical hours and generated a total of TB of output for the 3D MHD variables, including the high-cadence output just described plus occasional snapshots of the full numerical domain. The initial magnetic configuration, total duration, and total output for all cases are summarized in Table 1.
@@ 4 Artifacts in MHD Simulation and Solutions
These oscillations are barely detectable under a weak mean magnetic field; specifically, Case 5 is intermittently impacted, whereas Case 4 remains mostly unaffected. As a test, we also simulated a case with a stronger, G added magnetic field, which produced even stronger oscillations (Figure 24(b)). Efforts to stabilize this simulation by reducing the CFL parameter and adjusting velocity parameters were unsuccessful. A potential solution involves expanding the simulation box upwards and incorporating a realistic corona. This would significantly increase computational demands, exceeding the SPIn4D project’s scope. Consequently, we limited our simulations to a maximum of G magnetic field strength.
@@
We note that these aforementioned artifacts are not expected to impact our results, as illustrated by Figure 24. The upper panel (a) is for Case 5. The regions of numerical instability produce very high temperatures, exceeding K, saturating the color table in white. These regions produce both the strong oscillations and shocks in the surrounding plasma, but are basically confined to strong magnetic field regions. The magnitude of the oscillations decreases as the density rises in the lower regions of the atmosphere, such that the oscillations are largely confined to layers where . Given the distribution of the response function, as shown in Figure 8(a)–(d), the regions critical for the formation of Fe I lines remain unaffected.
@@ Appendix A Changes in SIR3D and Data Organizing
To enhance the handling of half-integer quantum levels for the Fe I nm lines, new atomic state symbols have been updated in the src/interface.f90 module of the 3D version of SIR. The orbital angular momentum range has been expanded from six integer levels to include 13 integer and 11 half-integer levels. The element abundance data was updated in both src/interface.f90 and src/leyendo.f. We have enhanced the selection of optical depth ranges for forward synthesis by introducing new parameters in synth/model.py: clip_tau, clip_tau_min, and clip_tau_max. To address precision issues in the upper layers of MURaM simulations that cause duplication at very small optical depths, line synthesis is now constrained to layers within . We modified the SIR3D code to support a new ’MURAM’ atmosphere that uses solar Cartesian coordinates to match the DL-NIRSP axis ordering and the conventions used in the Python package for the solar community (SunPy). MURaM’s SSD simulation coordinates (, , ) correspond to the SIR3D’s coordinates (, , ) and to solar Cartesian coordinates (, , ), thus a transpose(1,0,2) transformation is applied to all memmap objects in synth/multiprocessing.py to align with the solar Cartesian coordinate system. See our project website for a detailed explanation of these changes.
On the SPIn4D project website, aur datasets are stored within corresponding case directories, whose names are characterized by the mean magnetic field strength. The directories are labeled as follows: SPIN4D_SSD, SPIN4D_SSD_50G, SPIN4D_SSD_50G_V, SPIN4D_SSD_100G, SPIN4D_SSD_200G, and SPIN4D_SSD_Large, corresponding to Cases 1 through 6. For instance, the file subdomain_0.051405 in the directory SPIN4D_SSD_50G corresponds to the mass density data from the simulation output at time index "051405", where "0" is the variable index. Variable indices from 0 to 11 represent mass density (), velocities (, , ), internal energy (), magnetic fields (, , ), temperature (), pressure (), electron number density (), and optical depth (), the shift of the vector components reflecting the coordinate transformations between SIR and MURaM aforementioned. Furthermore, stokes-051405-6302.h5 corresponds to the Fe I 630 nm line at the corresponding output time index. Detailed guidance on accessing and interpreting the data is provided in our online tutorial555https://github.com/ifauh/spin4d-data.

### v3

@@ Abstract
The National Science Foundation’s Daniel K. Inouye Solar Telescope (DKIST) will provide high-resolution, multi-line spectropolarimetric observations that are poised to revolutionize our understanding of the Sun. Given the massive data volume, novel inference techniques are required to unlock its full potential. Here, we provide an overview of our “SPIn4D” project, which aims to develop deep convolutional neural networks (CNNs) for estimating the physical properties of the solar photosphere from DKIST spectropolarimetric observations. We describe the magnetohydrodynamic (MHD) modeling and the Stokes profile synthesis pipeline that produce the simulated output and input data, respectively. These data will be used to train a set of CNNs that can rapidly infer the four-dimensional MHD state vectors by exploiting the spatiotemporally coherent patterns in the Stokes profile time series. Specifically, our radiative MHD model simulates the small-scale dynamo actions that are prevalent in quiet-Sun and plage regions. Six cases with different mean magnetic fields have been conducted; each case covers six solar-hours, totaling TB in data volume. The simulation domain covers at least Mm with km spatial resolution, extending from the upper convection zone up to the temperature minimum region. The outputs are stored at a 40 s cadence. We forward model the Stokes profile of two sets of Fe I lines at and nm, which will be simultaneously observed by DKIST and can better constrain the parameter variations along the line of sight. The MHD model output and the synthetic Stokes profiles are publicly available, with 13.7 TB in the initial release.
@@ 1 Introduction
With the help of the above new facilities, spectropolarimetric observations (typically in the form of wavelength dependent Stokes profiles) for solar physics will significantly expand in the near future. These advanced observations are sensitive to the plasma state variables across the solar atmosphere (del Toro Iniesta & Ruiz Cobo, 2016; Bellot Rubio & Orozco Suárez, 2019, and reference therein), offering a unique tool to probe the solar environment in different heights through sophisticated inversion techniques, whereby the multi-dimensional state of the plasma is inferred from 2D maps of the polarized spectra (Ruiz Cobo & del Toro Iniesta, 1992; Asensio Ramos et al., 2008; Socas-Navarro et al., 2015; Milić & van Noort, 2018; delaCruzRodriguez et al., 2019; Quintero Noda et al., 2021, 2023; Ruiz Cobo et al., 2022; Asensio Ramos & Díaz Baso, 2019; Pastor Yabar et al., 2019; Borrero et al., 2019). For example, initial DKIST/ViSP data measurements of magnetic fields in both quiet and plage regions highlight the utility of spectropolarimetric observation in probing plasma dynamics in otherwise inaccessible environments (Campbell et al., 2023; da Silva Santos et al., 2023; Kuridze et al., 2024). Developing good inversion techniques is thus required to tackle a host of unanswered questions in both solar and plasma physics in general, for instance, probing the mechanisms behind local magnetic dynamos (Vögler & Schüssler, 2007; Stenflo, 2012; Rempel, 2014; Lord et al., 2014), understanding the flux of helicity and energy across the solar surface (Welsch, 2015; Schuck & Antiochos, 2019; Lumme et al., 2019; Thalmann et al., 2021; Liu & Schuck, 2012a; Liu et al., 2014, 2023), and unraveling the magnetic foundations necessary for solar eruptions and coronal heating (Antiochos, 1998; Antiochos et al., 1999; Moore et al., 2001; Priest & Forbes, 2002; Sun et al., 2013; Wang et al., 2015; Liu et al., 2016; Chitta et al., 2017; Wyper et al., 2017; Wang et al., 2017; Samanta et al., 2019).
@@
To address the challenges outlined above and take advantage of the potential of DL, we launched the “Spectropolarimetric Inversion in Four Dimensions with Deep Learning” (SPIn4D) project222https://ifauh.github.io/SPIN4D/. to train CNN models on radiative MHD (RMHD) simulations with a significantly larger dataset, i.e., larger domain size, longer evolution, and more cases with a variety of mean magnetic field strengths, compared to previous efforts. As discussed below, this provides more statistically independent snapshots useful for training. Detailed information about the data, including access methods, are publicly available333Several data access methods are offered for Data Release 1 at http://dtn-itc.ifa.hawaii.edu/spin4d/DR1/..
Figure 2 presents a schematic overview of key elements of each phase of the project. Our focus will be on photospheric regions with intermediate field strengths (between quiet sun and plage, or up to a few hundred Gauss when spatially averaged), characterized by relatively simple field geometries. These regions are expected to be prevalent in the initial years of DKIST’s operation. Simulations are carried out using the Max-Planck University-of-Chicago Radiative MHD code (MURaM, Vögler et al., 2005; Rempel et al., 2009; Cheung et al., 2010; Rempel, 2012, 2014) and take the quiet Sun small-scale dynamo simulation (SSD) of Rempel (2014) as the point of departure. The fact that these MURaM simulations use gray radiative transfer is not expected to impact the results here. As both the training and the evaluation steps use synthetic data, the ML models should simply learn the mapping between the self-consistent inputs and outputs, without regards to the detailed physics. When applying to real observations, however, we do expect ML models trained on more realistic, non-gray radiative transfer simulations (Vögler, 2004) to perform better, at the expense of increased computational costs.
Next, we employ forward-modeling of radiative transfer through these simulations using a new version of the SIR code (SIR3D) and the Departure Coefficient aided Stokes Inversion based on Response Functions (DeSIRe, Ruiz Cobo et al., 2022) code to synthesize Stokes profiles for multi-line observations. We have selected the well-studied Fe I lines at nm, nm, nm, and nm due to their significant Landé factors (, , , and , respectively) and their wide-spread use in ground and space-based observations with high Stokes SNR.
The third phase involves developing CNN models that aim to accurately correlate time series of observational data, especially those from the DL-NIRSP instrument at DKIST, with precise 4D MHD states. The models will be rigorously trained and evaluated using the data generated in the first two steps. The ability of this model will be compared with the SIR inversion code as a base line. Higher-level variables such the vector velocity and associated Poynting flux across the photosphere may be estimated either based on the inversion results as in conventional methods (e.g., Liu & Schuck, 2012b; Kazachenko et al., 2014), or ultimately directly from our ML models that encode the temporal information with a time-series of Stokes profiles as input.
@@ 2 Solar Atmosphere Simulation
All Cases 1–6 were simulated for about six physical hours (including the initial two hours of relaxation) and generated a total of TB of output for the 3D MHD variables, including the high-cadence output just described plus occasional snapshots of the full numerical domain. The initial magnetic configuration, total duration, and total output for all cases are summarized in Table 1.
@@ 4 Artifacts in MHD Simulation and Solutions
It is worth noting that such oscillations do not appear in Rempel (2014), and they are barely detectable for the cases with weaker mean magnetic field. Specifically, Case 5 is intermittently impacted, whereas Case 4 remains mostly unaffected. As a test, we also simulated a case with a stronger, G added magnetic field, which produced even stronger oscillations (Figure 24(b)). The problem arises from both stronger magnetic fields and the more extended vertical domain. The combination leads to large values exceeding the imposed limit. Efforts to stabilize this simulation by reducing the CFL parameter and adjusting velocity parameters were unsuccessful. A potential solution involves expanding the simulation box upwards and incorporating a realistic corona. This would significantly increase computational demands, exceeding the SPIn4D project’s scope. Consequently, we limited our simulations to a maximum of G magnetic field strength.
@@
We note that these aforementioned artifacts are not expected to impact our results, as illustrated by Figure 24. The upper panel (a) is for Case 5. The regions of numerical instability produce very high temperatures, exceeding K, saturating the color table in white. These regions produce both the strong oscillations and shocks in the surrounding plasma, but are basically confined to strong magnetic field regions. Since the MURaM code uses a numerical diffusivity scheme that can significantly enhance diffusivities in regions where monotonicity changes (see Sect. 2 in Rempel, 2014) where needed and keep them low everywhere else. This sufficiently decays and prevents the spread of these oscillations to the lower and higher atmosphere, mainly trapped around Mm where they are generated. The magnitude of the oscillations decreases as the density rises in the lower regions of the atmosphere, such that the oscillations are largely confined to layers where . Given the distribution of the response function, as shown in Figure 8(a)–(d), the regions critical for the formation of Fe I lines remain unaffected.
@@ Appendix A Changes in SIR3D and Data Organizing
To enhance the handling of half-integer quantum levels for the Fe I nm lines, new atomic state symbols have been updated in the src/interface.f90 module of the 3D version of SIR. The orbital angular momentum range has been expanded from six integer levels to include 13 integer and 11 half-integer levels. The element abundance data was updated in both src/interface.f90 and src/leyendo.f. We have enhanced the selection of optical depth ranges for forward synthesis by introducing new parameters in synth/model.py: clip_tau, clip_tau_min, and clip_tau_max. To address precision issues in the upper layers of MURaM simulations that cause duplication at very small optical depths, line synthesis is now constrained to layers within . We modified the SIR3D code to support a new ‘MURAM’ atmosphere that uses solar Cartesian coordinates to match the DL-NIRSP axis ordering and the conventions used in the Python package for the solar community (SunPy). MURaM’s SSD simulation coordinates (, , ) correspond to the SIR3D’s coordinates (, , ) and to solar Cartesian coordinates (, , ), thus a transpose(1,0,2) transformation is applied to all memmap objects in synth/multiprocessing.py to align with the solar Cartesian coordinate system. See our project website for a detailed explanation of these changes.
On the SPIn4D project website, our datasets are stored within corresponding case directories, whose names are characterized by the mean magnetic field strength. The directories are labeled as follows: SPIN4D_SSD, SPIN4D_SSD_50G, SPIN4D_SSD_50G_V, SPIN4D_SSD_100G, SPIN4D_SSD_200G, and SPIN4D_SSD_Large, corresponding to Cases 1 through 6. For instance, the file subdomain_0.051405 in the directory SPIN4D_SSD_50G corresponds to the mass density data from the simulation output at time index "051405", where "0" is the variable index. Variable indices from 0 to 11 represent mass density (), velocities (, , ), internal energy (), magnetic fields (, , ), temperature (), pressure (), electron number density (), and optical depth (), the shift of the vector components reflecting the coordinate transformations between SIR and MURaM aforementioned. Furthermore, stokes-051405-6302.h5 corresponds to the Fe I 630 nm line at the corresponding output time index. Detailed guidance on accessing and interpreting the data is provided in our online tutorial555https://github.com/ifauh/spin4d-data. It is also available on Zenodo https://doi.org/10.5281/zenodo.13879854 (catalog doi:10.5281/zenodo.13879854).